OpenStudy (anonymous):

Meg has a can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of orange juice and water in the can. Define the variable used in the equation. Part B: How many liters of orange juice are present in the can? Show your work.

OpenStudy (anonymous):

0.2 V = 1 liter is a start, where V = volume of can

OpenStudy (texaschic101):

If orange juice is 80%, water is the other 20% and water = 1 liter, so 20% of the can is 1 liter. let x represent total amount of liters in the can 1/20 = x/100 -- (1 liter to 20% = x liters to 100%) cross multiply (20)(x) = (1)(100) 20x = 100 -- divide by 20 x = 5 total amount of liters in the can is 5. check.. water...20%.....20% of 5 = .20(5) = 1 liter orange juice ...80%....80% of 5 = .80(5) = 4 liters total = 100%....4 liters + 1 liters = 5 liters